Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.
7x + 9y = -27
-2x - 7y= 21

Answer by richwmiller(17219)


7x + 9y = -27
-2x - 7y= 21
multiply first by 2
second by 7
2*7x + 2*9y = -2*27
-2*7x - 7*7y= 147
-14x-49y=147
14x+18y=-54
-31y=93
y=-3
-2x-7y=21
-2x-7*(-3)=21
-2x=0
x=0
x=0 y=-3